some oxide superconductors adopt a crystal structure similar to that of perovskite CaTiO3. the unit cell is cubic with a Ti4+ ion in each corner, a Ca2+ ion in the body center and O2- ions ate the midpoint of each edge.

(a) if the unit cell edge length is 3.84 A, what is the density of perovskite in g/cm3??
Answer #1

a = 4r/\sqrt{}3 for BCC

a = 4*3.84*10-8/1.732 = 8.87*10-8 cm

d    = ZM/N0a3

      = 2*136/6.023*1023 *(8.87*10-7)3

     = 272/53.424*102

     = 0.0509g/cm3 >>>>> answer